please give detailed process and explaination
of r RV-300 Ω, and an of r
Answer #1

a) Rp = 75 * 300 / 375 = 60 ohms

60 + 3 + 100 = 163 ohm

I = 18 / 163

ammeter reading = 0.1104 A

b) v = 0.1104 * 60 = 6.625 V

voltmeter reading = 6.625 V

c) resistance = 6.625 / 0.1104

= 60.00 ohm